With current technology, suppose a firm is producing 500 shirts daily. Also assume that the least-cost combination of resources in producing those shirts is 10 units of labor, 8 units of land, 3 units of capital, and 4 units of entrepreneurial ability, selling at prices of $50, $40, $60, and $30, respectively. If the firm can sell these 500 shirts at $2 per unit, what is its total revenue? Its total cost? Its profit or loss? Will it continue to produce shirts? If this firm’s situation is typical for the other makers of shirts, will resources flow toward or away from this apparel?


1
Expert's answer
2021-03-08T09:21:31-0500

The total revenue after selling the shirts


= $2×500 = $1000.


The total cost of making the shirts = (10×50) +(8×40) +(3×60) +(4×30) = 500+320+180+120


=$1120.


The total cost here is more than its total revenue. So, the firm will run in a loss.


It will not continue to produce shirts as long as the price if shirts per unit does not rise which could have eventually overcome the cost and incurred profit.If the price of each shirt remains to be $2 per unit, and the cost of production remains to be the same, firm will not further continue to produce shirts as this will continuously incur loss.


Resources will not flow toward this apparel since the firm is making losses, hence other firms like the ones selling the factors of production to the shirt firm will tend to shift their allocation of resources to other firms which are profit maximizers.